OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3) Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3) @Mathematics

OpenStudy (anonymous):

4x - 24 = -2x + 6 6x = 30 x = 5

OpenStudy (agreene):

start by distributing 4 and -2 then combine x's on one side coeffs on the other divide by a where ax x=5

OpenStudy (anonymous):

4x-24=-2x+6 4x=-2x+30 6x=30 x=5

OpenStudy (anonymous):

4(x - 6) = -2(x - 3) 4x-24 = -2x+6 4x+2x = 6+24 6x = 30 x = 5